OpenStudy (anonymous):

Which of the following circle equations represents a circle whose center is at (0, 0)?

OpenStudy (anonymous):

x2 + y2 = 45 (x − 3)2 + (y − 5)2 = 0 (x + 3)2 + (y − 3)2 = 0 x2 + (y − 1)2 = 1

OpenStudy (anonymous):

we can eassily say that first eqution... that is x2 + y2 = 45

OpenStudy (anonymous):

becouse circle eqution is (x - a)^2 + (y - b)^2 = 0 hear (a,b) are the center point..... compare the eqution with the give you will get the answer....